A small country produces sugar domestically but also imports from other countries. The domestic supply curve is Qsdomestic=5p and the foreign supply curve is Qsforeign=10p . The demand of sugar in this country is given by Qd=100−5p .

Give the coefficients of the aggregate (i.e. combined) supply curve for this market, expressed as Qsaggregate=a+bp ..what is the value of a and b..what is the equilibrium price and quantity imported
1
Expert's answer
2020-06-09T16:49:52-0400
"Qs_{aggregate}=5p+10p=15p"

a=0, b=15


"100-5p=10p"


"15p=100"


"p=6\frac{2}{3}"